# Task-Adaptive Pretrained Language Models via Clustered-Importance Sampling

### v1

@@ 5.1 Clustering
The representation is important: the examples in a cluster are close in the embedding space. Our independence assumption, Equation 1, assumes that the loss in a cluster $c$ is the same regardless whether its data originates from ${D}^{\rm g}$ or ${D}^{\rm s}$, i.e.
@@
In practice, it is sufficient that the embedding space reflects the similarity of the loss gradient, i.e. if the gradients of the loss over a generalist cluster ${D}^{\rm g}\cap K(c)$ is correlated with the gradient over a specialist cluster ${D}^{\rm s}\cap K(c);\theta)$, the model trained on the former improves on the later. Figure 7 shows that the SBERT representation is more adapted and yields better results than LSI for all settings.
@@
The number of clusters is a trade-off and its optimum is at 260k for most of our experiments. There are multiple factors at play when the number of clusters varies. A smaller number of clusters implies larger clusters: our hypothesis, Equation 3, is then stronger as it assumes loss similarity on large areas of the embedding space. At the limit, with one cluster, this hypothesis assumes that the specialist loss and generalist loss are identical everywhere. Conversely, as the number of clusters gets larger, the estimation of the cluster density on the small specialist set $P(c|{\cal D}^{\rm s})\simeq P(c|{D}^{\rm s})$ gets less accurate. The estimator risks overfitting, i.e. favoring clusters frequent in the training set ${D}^{\rm s}$ but not as frequent on other samples from ${\cal D}^{\rm s}$. Increasing the number of clusters also risks reducing the effective training set size: the specialist data could be mostly concentrated in a few clusters, corresponding to a small fraction of the overall generalist set ${D}^{\rm g}$.
We explore these aspects on MMLU. We first measure the number of repeated examples when training models with CRISP pretraining for different number of clusters. Figure 7 shows the number of repetitions for each quantile of the training set. Even for 16.7m clusters, only a small minority of training examples are repeated beyond 10 times and the average number of occurrences of the training is examples 1.94, well within commonly recommended values (Muennighoff et al., 2023; Xue et al., 2023). Even if exact repetitions do not account for the poorer performance of the 16.7m cluster setting, its training set might be less diverse and the model might generalize well only in a small neighborhood of its training set. We plot Figure 7 to evaluate if the Redpj2 examples with good perplexity concentrate around $D^{\rm s}$, the MMLU training set. This plot shows that the benefit of CRISP over base is indeed correlated with the distance to $D^{\rm s}$. As the number of cluster increases to 16.7m, the benefit over base concentrate in an area with very few samples. For comparison, we plot the 2 middle quartiles [0.25, 0.75] where most of the MMLU test data concentrate in gray. We remark that MMLU test data mostly lies in an area where the perplexity of IS 16.7m is low.
@@
Figure 8 shows the perplexity during training for CRISP runs on MMLU. On Figure 8(a), the perplexity is computed from the reweighed loss on Redpj2. This is the loss optimized during pretraining. It shows that when the number of cluster increases the sampled training set is less diverse and corresponds to an easier learning problem ($<5$ PPL). On Figure 8(b), the perplexity is computed on the MMLU data itself, on the training set (plain) and on the test set (dotted). The scale of both plot is different: the resampled perplexities on Redpj2 are therefore not a good approximation of the MMLU perplexities. This quantifies the error resulting from our assumption, Equation 3. We also remark overfitting for 16.7m clusters, the only case with better MMLU perplexity for train than for test. Finally, we notice that the gray area in Figure 7 fails to show that 260k cluster would have the best perplexity, which highlights that SBERT distance to the training data is not the only factor explaining model performance.
@@ 5.3 Different Amount of Training Data
This section varies both (i) the amount of generalist data available to sample the CRISP dataset from and (ii) the amount of specialist data for inferring the CRISP weights.
When specialist data concentrates on a few clusters, CRISP often samples generalist data from the same clusters, which can be problematic when the generalist set is small. We restrict the pretraining set to 700B and 120B tokens (downsampling Redpj2 by resp. $\sim$ 50x and $\sim$300x). Our pretraining runs use 120B tokens, so a base run never repeats in all settings. When CRISP is applied, some tokens are repeated. Figure 11(a) shows that, when restricting to 120B tokens, the number of repetition becomes high (22.5 on average) and CRISP is ineffective after 256k steps.
@@ 5.4 Task-Transfer and Multitasking
We perform cross-task evaluation, i.e. targeting a task A and evaluating on a task B, we also pretrain a multitask models with CRISP averaged weights from multiple tasks. Our results for the 1.3B models are in Table 1, we also report cross-task evaluation results for different model sizes in Appendix H. Cross-task evaluations show that, perhaps unsurprisingly, the best results on a task A are obtained when pretraining for task A. Transfer differs across tasks: CRISP targeting MMLU gives better results than base for all tasks, which is not the case for CRISP targeting ARC or RWDB-R. The multi-task result which mixes the histograms with the same weight (1/3 for ARC, MMLU and RWDB-R) gives the best result on averaged multitask accuracy. Surprisingly, on RWDB-R, this setting slightly outperforms targeting RWDB-R itself.
@@ 6 Conclusions
A specialist LM is interesting since a small specialist LM can outperform a larger generalist LM on its targeted domain while having a lower inference cost. We explore pretraining specialist LMs when little specialization data is available, a common setting that prevents pretraining of dedicated LMs. We evaluate different method that modifies the distribution of a generic training set guided by little specialist data. Our experiments highlight the benefit of clustered importance sampling: i.e. resampling the generic set such that its cluster histogram matches the specialist data. Our findings show that pretraining with this method provides strong models both for LM and question answering tasks. These benefits are studied across model size, training set size and clustering methods. We also explore ways to lower the training cost of specialist models addressing by showing their benefit on shorter training runs, continued pretraining and multitask settings.
Our work shows that a simple, scalable importance sampling method can provide effective specialist LMs, even from little specialization data. Since clustered importance sampling is modality-agnostic, we foresee extensions of this work to other modalities, including vision and audio.
Acknowledgements
We thank Matteo Pagliardini and Angelos Karathopoulos for their help throughout the project. We also thank Maartje ter Hoeve, Yizhe Zhang and Navdeep Jaitly for their suggestions and comments.
@@ Appendix A Scalable Clustering
We cluster the Redpj2 dataset with hierarchical clustering. We build a clustering tree. Each node in the tree is associated with a cluster centroid. The examples traverse the tree from top to bottom, selecting the node corresponding to the closest centroids among the current node’s children.
@@
Appendix B Dataset Statistics
@@
Appendix C Architectures & Hyperparameters
@@
Appendix D MCQ Evaluation
For the multiple-choice-question tasks, we use the LM eval harness (Gao et al., 2024). For each task, the evaluated model estimates the (log) probablity of each answer $a$ given the context $c$, i.e. $\log P(a|c)$. The question contains the task prompt concatenated with the current question, while the answer contains the answer text. With this strategy, the model does not have access to all proposed answer choices in the prompt. Table 7 reports our prompt. For all evaluation, we use the above prompt without example questions, i.e. a zero-shot evaluation (Brown et al., 2020). Accuracy is computed by verifying if the highest score is assigned to the correct answer. The scores correspond to log probabilities for ARC-E and RWDB-R, while ARC-C, MMLU uses normalized scores, i.e. log probabilities divided by the number of characters in the answer.
@@ Appendix D MCQ Evaluation
Appendix E Supplementary Results for Language Modeling Tasks
We measure the fraction of examples where the pretrained model is better than the base model. We measure this rate both on held-out data from ${\cal D}^{\rm g}$ (measured on the 360m tokens from Redpj2 valid) and on held-out data from ${\cal D}^{\rm g}$ (measured on the full Pile validation set). The results in Table 8 shows that the model trained with importance sampling improves perplexity on most specialist documents (right column). Its training on the importance sampled distribution utilize model capacity mostly on data close to the domain of interest, this relieves the model from fitting well most of the generic data, hence most generic documents have higher perplexity with CRISP (left column).
For completeness, we also report the perplexity numbers from Figure 4 in Table E.
@@ Appendix E Supplementary Results for Language Modeling Tasks
Appendix F Supplementary Results for Multiple Choice Questions
@@
Appendix G Comparing the results of DoGE and Importance Sampling
@@
Appendix H Task Transfer for 350m, 1.3B and 7B models

### v2

@@ 5.1 Clustering
The representation is important: examples in the same cluster are close in the embedding space. Our independence assumption, Equation 1, assumes that the loss in a cluster $c$ is the same regardless whether its data originates from ${D}^{\rm g}$ or ${D}^{\rm s}$, i.e.
@@
In practice, it is sufficient that the embedding space reflects the similarity of the loss gradient, i.e. if the gradients of the loss over a generalist cluster ${D}^{\rm g}\cap K(c)$ is correlated with the gradient over a specialist cluster ${D}^{\rm s}\cap K(c);\theta)$, the model trained on the former improves on the later. Figure 7 shows that the SBERT representation yields better results than LSI for all settings.
@@
The number of clusters is a trade-off, and its optimum is at 260k for most of our experiments. There are multiple factors at play when the number of clusters varies. A smaller number of clusters implies larger clusters: our hypothesis, Equation 3, is then stronger, as it assumes loss similarity on large areas of the embedding space. At the limit, with one cluster, this hypothesis assumes that the specialist loss and generalist loss are identical everywhere. Conversely, as the number of clusters gets larger, the estimation of the cluster density on the small specialist set $P(c|{\cal D}^{\rm s})\simeq P(c|{D}^{\rm s})$ gets less accurate. The estimator risks overfitting, i.e. favoring clusters frequent in the training set ${D}^{\rm s}$ but not as frequent on other samples from ${\cal D}^{\rm s}$. Increasing the number of clusters also risks reducing the effective training set size: the specialist data could be mostly concentrated in a few clusters, corresponding to a small fraction of the overall generalist set ${D}^{\rm g}$.
@@
We explore these aspects on MMLU. We first measure the number of repeated examples when training models with CRISP pretraining for different number of clusters. Figure 7 shows the number of repetitions for each quantile of the training set. Even for 16.7m clusters, only a small minority of training examples are repeated beyond 10 times and the average number of occurrences of the training is examples 1.95, well within commonly recommended values (Muennighoff et al., 2023a; Xue et al., 2023). Even if exact repetitions do not account for the poorer performance of the 16.7m cluster setting, its training set might be less diverse and the model might generalize well only in a small neighborhood of its training set. We evaluate if the Redpj2 examples with good perplexity concentrate around $D^{\rm s}$, the MMLU training set. Figure 7 shows that the benefit of CRISP over base is indeed correlated with the distance to $D^{\rm s}$. As the number of cluster increases to 16.7m, the benefit over base concentrate in an area with very few samples. For comparison, we plot the 2 middle quartiles [0.25, 0.75] where most of the MMLU test data concentrate in gray. We remark that MMLU test data mostly lies in an area where the perplexity of IS 16.7m is low.
Figure 8 shows the perplexity for CRISP runs on MMLU. On Figure 8(a), the perplexity is computed from the reweighed loss on Redpj2. This is the loss optimized during pretraining. It shows that when the number of cluster increases the sampled training set is less diverse and corresponds to an easier learning problem ($<5$ PPL). On Figure 8(b), the perplexity is computed on the MMLU data itself, on the training set (plain) and on the test set (dotted). The scale of both plot is different: the resampled perplexities on Redpj2 are therefore not a good approximation of the MMLU perplexities. This quantifies the error resulting from our assumption, Equation 3. We also see overfitting for 16.7m clusters, the only case with better MMLU perplexity for train than for test. Finally, we notice that the gray area in Figure 7 fails to show that 260k cluster would have the best perplexity, which shows that SBERT distance to the training data is not the only factor explaining model performance.
@@ 5.3 Different Amount of Training Data
This section varies both the amount of generalist data available to sample the CRISP dataset from and the amount of specialist data for inferring the CRISP weights. When specialist data concentrates on a few clusters, CRISP often samples generalist data from the same clusters, which can be problematic when the generalist set is small. We restrict the pretraining set to 700B and 120B tokens (downsampling Redpj2 by resp. $\sim$ 50x and $\sim$300x). Our pretraining runs use 120B tokens, so a base run never repeats in all settings. When CRISP is applied, some tokens are repeated. Figure 11(a) shows that, when restricting to 120B tokens, the number of repetition becomes high (22.5 on average) and CRISP is ineffective after 256k steps.
@@ 5.4 Task-Transfer and Multitasking
We perform cross-task evaluation, i.e. targeting a task A and evaluating on a task B, we also pretrain a multitask models with CRISP averaged weights from multiple tasks. Our results for the 1.3B models are in Table 1, we also report cross-task evaluation results for different model sizes in Appendix J. Cross-task evaluations show that, perhaps unsurprisingly, the best results on a task A are obtained when pretraining for task A. Transfer differs across tasks: CRISP targeting MMLU gives better results than base for all tasks, which is not the case for CRISP targeting ARC or RWDB-R. The multi-task result which mixes the histograms with the same weight (1/3 for ARC, MMLU and RWDB-R) gives the best result on averaged multitask accuracy. Surprisingly, on RWDB-R, this setting slightly outperforms targeting RWDB-R itself.
@@ 6 Conclusions
A small specialist LM is interesting since it can outperform a larger generalist LM on its targeted domain while having a lower inference cost. We explore pretraining specialist LMs when little specialization data is available, a common setting that prevents pretraining of dedicated LMs. We evaluate different methods that modify the distribution of a generic training set guided by little specialist data. Our experiments highlight the benefit of clustered importance sampling: i.e. resampling the generic set such that its cluster histogram matches the specialist data. Our findings show that pretraining with this method provides strong models both for LM and question answering tasks. We also explore ways to lower the training cost of specialist models by showing their benefit on shorter training runs, continued pretraining and multitask settings. Our work shows that a simple, scalable importance sampling method can provide effective specialist LMs, even from little specialization data. Since clustered importance sampling is modality-agnostic, we foresee extensions of this work to other modalities, including vision and audio.
Acknowledgments
We thank Angelos Katharopoulos, Matteo Pagliardini and Anastasiia Filippova for their advice throughout this project. We thank the anonymous reviewers for their suggestions and comments.
@@ Appendix A Scalable Clustering
We cluster the generic dataset (Redpj2) with hierarchical clustering. We build a clustering tree. Each node in the tree is associated with a cluster centroid. The examples traverse the tree from top to bottom, selecting the node corresponding to the closest centroids among the current node’s children.
@@
Appendix B Clustering & Embedding Cost
The computational cost of k-means clustering is negligible compared to the cost of computing text embeddings. Most of the experiments in this work are performed with SBERT MiniLM-L6-v2 Reimers & Gurevych (2019). Embedding the 34.6T tokens of Redpj2 amounts to $5.4$k GPU hours on a reference NVidia H100 GPU. Other models can provide better embeddings at a higher cost. We examined the results of clustering accuracy, MTEB evaluation Muennighoff et al. (2023b), versus embedding cost (for Redpj2 in GPU hours on H100) in Table 2.
Table 2: Embedding Cost versus Clustering Accuracy.
Clustering method Cost (GPUh) Accuracy (%)
all-MiniLM-L6-v2 Reimers & Gurevych (2019) 5.4k 41.94
e5-large-v2 Wang et al. (2022) 91.4k 44.26
e5-base-v2 Wang et al. (2022) 27.4k 44.10
all-mpnet-base-v2 Reimers & Gurevych (2019) 28.6k 43.69
gte-base-v1.5 Li et al. (2023) 87.4k 47.90
gte-small Li et al. (2023) 13.3k 44.89
In this cost-benefit table, gte-small stands out. We clustered the Redpj2 dataset with embedding from this model and we report MCQ results with CRIPS over this clustering. We compare these results with LSI and SBERT from Figure 7 in the main text. The results in Table 3 show that these embeddings are beneficial, especially with a small number of clusters.
Table 3: MCQ accuracy with CRISP for different embedding methods.
Clusters Emb. Arc-E Arc-C MMLU
64 LSI 65.0 31.8 31.1
SBERT 65.2 33.0 31.2
GTE 67.8 33.8 31.5
4096 LSI 67.6 35.4 31.6
SBERT 69.7 36.5 32.6
GTE 69.9 37.0 32.8
262k LSI 66.5 36.4 31.3
SBERT 71.2 38.6 33.4
GTE 69.3 37.6 33.4
16m LSI 53.8 29.6 29.0
SBERT 62.3 33.8 30.8
GTE N/A N/A 31.7
Appendix C Dataset Statistics
@@
Appendix D Architectures & Hyperparameters
@@
Appendix E MCQ Evaluation
For multiple-choice questions, we use the LM eval harness (Gao et al., 2024). For each task, the evaluated model estimates the (log) probability of each answer $a$ given the context $c$, that is, $\log P(a|c)$. The question contains the task prompt concatenated with the current question, while the answer contains the answer text. With this strategy, the model has no access to alternative answer choices proposed in the prompt. Table 9 reports our prompt. For all evaluations, we use the above prompt without example questions, that is, a zero-shot evaluation (Brown et al., 2020). Accuracy is calculated by verifying whether the highest score is assigned to the correct answer. The scores correspond to log probabilities for ARC-E and RWDB-R, while ARC-C, MMLU uses normalized scores, i.e. log probabilities divided by the number of characters in the answer.
@@ Appendix E MCQ Evaluation
Appendix F Supplementary Results for Language Modeling Tasks
We measure the fraction of examples where the pre-trained model is better than the base model. We measure this rate both on the held-out data from ${\cal D}^{\rm g}$ (measured on the 360m tokens from Redpj2 valid) and on held-out data from ${\cal D}^{\rm g}$ (measured on the full Pile validation set). The results in Table 10 show that the model trained with importance sampling improves perplexity on most specialist documents (right column). Its training on the importance sampled distribution utilize model capacity mostly on data close to the domain of interest, this relieves the model from fitting well most of the generic data, and hence most generic documents have higher perplexity with CRISP (left column).
For completeness, we also report the perplexity numbers of Figure 4 in Table F.
@@ Appendix F Supplementary Results for Language Modeling Tasks
Appendix G Supplementary Results for Multiple Choice Questions
@@
Appendix H Comparing the results of DoGE and Importance Sampling
@@
Table 14: Comparison with Cross-Entropy Difference for MCQ Accuracy (%), 1.3B model.
Arc-E Arc-C MMLU
Base 58.4 27.4 30.0
CED 58.9 30.5 31.1
Doge 66.2 33.3 31.0
Classifier 68.2 36.7 32.4
CRISP 71.2 38.6 33.4
Appendix I Comparing CRISP and Cross-Entropy Difference (CED)
Contrasting the scores of two LMs (generalist and specialist) is a popular method for data selection (Moore & Lewis, 2010; Axelrod et al., 2011; Wang et al., 2018; Junczys-Dowmunt, 2018). We considered this method based on Wang et al. (2018): we obtain the specialist LM by fine-tuning a generalist LM on the specialist training set. We rely on a 350m parameter model for the selection. One should note that this method is particularly expensive since it requires scoring the entire Redpj2 dataset twice with an LM, which is more expensive than embedding and clustering the dataset. Our results show that CED improves over the classifier method but CRISP is significantly better, see Table 14.
Appendix J Task Transfer for 350m, 1.3B and 7B models